In the equation ( NO2 -- N2O5 + NO) 0.050 mol/l of NO2 exists at equilibrium what is the concentration of N2O5 at equilibrium


1
Expert's answer
2021-01-07T05:30:04-0500

Balance the equation

3NO2 "\to" N2O5 + NO

The Formula for Kc = [N2O5][NO]/[NO]3

We know that Kc is constant for a given reaction and for the given equation

[N2O5] = [NO]

"\implies" concentration of N2O5 ="\sqrt{[NO2]^3 *Kc}" = "\sqrt{(0.05)^3 * Kc}"